OpenStudy (anonymous):

@143007 Solve this 3x+24(25+32)=22 Solve for x

OpenStudy (callisto):

Assume you've no typing mistakes there 3x+24(25+32)=22 3x+24(57)=22 3x + 1368 =22 3x = 22-1368 Can you solve from here?